OpenStudy (anonymous):

y=2x^2-4x+12 in vertex form

OpenStudy (anonymous):

y=2x^2-4x+12 y/2=x^2-2x+6 y/2=(x^2-2x+1)+5 y/2=(x-1)^2+5